How many gallons of 40% antifreeze should be mixed with 10 gallons of 85% antifreeze to obtain 65% antifreeze mixture?
Answer by stanbon(75887)

How many gallons of 40% antifreeze should be mixed with 10 gallons of 85% antifreeze to obtain 65% antifreeze mixture?
------
Equation:
alcohol + alcohol = alcohol
0.40x + 0.85*10 = 0.65(x+10)
----
40x + 85*10 = 65x + 65*10
20*10 = 15x
-----
x = (2/3)20
---
x = 13 1/3 gallons (amt. of 40% solution needed)
=======================
Cheers,
Stan H.
=======================